OpenStudy (anonymous):

(448÷〖(-2)〗^3÷2×(-3)+1)/5

OpenStudy (anonymous):

This is when the equation editor comes in handy.

OpenStudy (anonymous):

what is that

OpenStudy (anonymous):

can you help me solve this

OpenStudy (anonymous):

\[\frac{ (\frac{ 448 }{ (-2)^3 }) }{ 2 } \times \frac{ -3+1 }{ 5 }\] ?

OpenStudy (anonymous):

(448÷〖(-2)〗^3÷2×(-3)+1 --------------------------------- 5

OpenStudy (anonymous):

(-2)^3 = -8 448/-8 = -56 -56/2=-28 -28 x -3 = 84 84+1 = 85 85/5 = 17 Do it in that order (BEDMAS/PEDMAS/BODMAS/PODMAS)

OpenStudy (anonymous):

is 17 the answer

OpenStudy (anonymous):

is that how you solve this problem